OpenStudy (anonymous):

31.5 grams of an unknown substance is heated to 102.4 degrees Celsius and then placed into a calorimeter containing 103.5 grams of water at 24.5 degrees Celsius. If the final temperature reached in the calorimeter is 32.5 degrees Celsius, what is the specific heat of the unknown substance? Show or explain the work needed to solve this problem, and remember that the specific heat capacity of water is 4.18 J/(°C x g)

OpenStudy (festinger):

By conservation of energy: \[Heat\:gained\:by\:water=Heat\:lost\:by\:unknown\:substance\]\[4.18*103.5*(32.5-24.5)=c(31.5)(102.4-32.5)\] Solve for c to get the specific heat capacity of the unknown substance in J/(°C g) If you don't know where the equation came from just ask.

OpenStudy (anonymous):

would you please explain

OpenStudy (festinger):

heat gained or lost in a substance is simply: mass x specific heat capacity x change in temperature This is because as we have more stuff to heat up, we need more heat. But different substances heat up differently, so there is a specific heat capacity to distinguish them and describe how easy it is to heat them or cool them.

OpenStudy (anonymous):

thanks. you're exetremely helpful

OpenStudy (anonymous):

what is c suppose to equal?

OpenStudy (anonymous):

@Festinger

OpenStudy (festinger):

1.57J/(g K)

OpenStudy (anonymous):

i didn't mean the answer lol. i meant like it stands for

OpenStudy (festinger):

read my first answer again: c = specific heat capacity of the unknown substance in J/(°C g) it is the measure of how much heat you need to give to the substance to bring 1g of it up by 1 degrees celcius.

OpenStudy (anonymous):

oh ok cool

OpenStudy (anonymous):

4.18* 103.5* (32.5-27.5) = c(31.5)(102.4-32.5) 432.63(5)=c(31.5)(69.9) 2163.15=c2201.85 c=.982 that's what i got...

OpenStudy (anonymous):

@Festinger

OpenStudy (festinger):

What is the initial temperature of water? is it 24.5C or 27.5C?

OpenStudy (anonymous):

24.5. don't you need to convert it to kelvin though

OpenStudy (festinger):

But your first line says: 4.18* 103.5* (32.5-27.5) = c(31.5)(102.4-32.5) Which says 27.5

OpenStudy (anonymous):

that's where i made my mistake